# Runtime Safety through Adaptive Shielding: From Hidden Parameter Inference to Provable Guarantees

### v1

@@ 4 Approach
Our approach has three main components. First, we introduce a novel safety-regularized objective. This objective is used during optimization and encourages the policy to converge toward a zero-violation policy. Second, we use a function encoder to represent underlying dynamics $T_{\theta}$, enabling online adaptation. Finally, we leverage this dynamics representation to construct an adaptive shield. The shield adjusts safe regions by conformal prediction and blocks unsafe actions online.
4.1 Safety-Regularized Objective
@@
To promote safe policy learning, we introduce a safety measure, $Q^{\pi}_{\text{safe}}(s,a,\theta)$, which quantifies the relative safety of action $a$ sampled from the policy $\pi$ in state $s$ under parameter $\theta$. Higher values of $Q^{\pi}_{\text{safe}}$ indicate actions with lower long-term costs under policy $\pi$. Since we aim to minimize the cost action-value function $Q^{\pi}_{C}$, higher $Q^{\pi}_{\text{safe}}$ values correspond to lower $Q^{\pi}_{C}$ values. Based on this intuition, we define $Q^{\pi}_{\text{safe}}(s,a,\theta)$ for an action $a$ sampled by $\pi(\cdot\mid s,b)$ as:
$Q^{\pi}_{\text{safe}}(s,a,\theta)=1-\frac{\pi(a\mid s,b)Q^{\pi}_{C}(s,a,\theta)}{V^{\pi}_{C}(s,\theta)+\epsilon}$ (6)
@@
where $b\in\mathbb{R}^{k}$ is a learned representation of $T_{\theta}$, and $\epsilon>0$ is a small constant for numerical stability.
This formulation bounds the value in $(0,1]$ by its design. A value of $1$ for $Q^{\pi}_{\text{safe}}(s,a,\theta)$ indicates two scenarios: 1) Safety: The policy selects an action $a$ resulting in zero long-term cost violations, i.e., $Q^{\pi}_{C}(s,a,\theta)=0$; or 2) Exploration: The probability of selecting action $a$ is small. In contrast, a value near 00 for $Q^{\pi}_{\text{safe}}(s,a,\theta)$ indicates that action $a$ substantially contributes to the expected cumulative cost $V^{\pi}_{C}(s,\theta)$, posing a higher risk compared to other action choices at state $s$. See Appendix G for details on the design choice.
To integrate this safety measure into policy optimization, we define an augmented action-value function, $Q^{\pi}_{\text{aug}}(s,a,\theta)=Q_{R}^{\pi}(s,a,\theta)+\alpha Q_{\text{safe}}^{\pi}(s,a,\theta)$, where $Q_{R}^{\pi}(s,a,\theta)$ is the reward action-value function, and $\alpha\geq 0$ is a hyperparameter balancing safety and reward. Our safety-regularized objective (SRO) is:
@@
$J_{\text{safe}}(\pi)=\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta),a_{0}\sim\pi(\cdot\mid s_{0},b)}\left[Q^{\pi}_{\text{aug}}(s_{0},a_{0},\theta)\right].$ (7)
@@
A larger $\alpha$ encourages the policy to prioritize safe actions that result in zero-violation costs or to select under-explored actions with lower assigned probabilities. Next, we introduce a proposition which justifies this choice of objective.
Proposition 1.
@@
Let $\Pi_{\text{zero-violation}}$ denote the set of zero-violation policies, defined as $\{\pi\mid J_{C}(\pi)=0\}$. Then, for any $\alpha\geq 0$, the optimal policy obtained by maximizing the safety-regularized objective function $J_{\text{safe}}(\pi)$ within $\Pi_{\text{zero-violation}}$ is equivalent to the optimal policy obtained by maximizing the standard reward objective $J_{R}(\pi)=\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta),a_{0}\sim\pi(\cdot\mid s_{0},b)}[Q_{R}^{\pi}(s_{0},a_{0},\theta)]$ within the same set of policies.
@@
For any policy within the zero-violation set, all actions sampled from the policy lead to $Q_{C}^{\pi}(s,a,\theta)=0$. By our design of the safety term, this condition implies $Q_{\text{safe}}^{\pi}(s,a,\theta)=1$. Substituting this into our regularized objective, $J_{\text{safe}}(\pi)$ simplifies to $J_{R}(\pi)+\alpha$. Therefore, for zero-violation policies, maximizing $J_{\text{safe }}(\pi)$ is equivalent to maximizing the standard reward objective $J_{R}(\pi)$, as $\alpha$ is a constant. ∎
Proposition 1 is significant because it proves that the safety regularization does not degrade performance unnecessarily when an agent already behaves safely. Specifically, it guarantees that if we focus only on the set of policies that satisfy all safety constraints, maximizing the safety-regularized objective is equivalent to maximizing the standard reward objective.
@@ 4.2 Inferring Hidden Parameters Online
To infer the underlying dynamics $T_{\theta}$ and predict the next state $s_{t+1}$ based on transition samples and $(s_{t},a_{t})$, we use a function encoder, denoted by $\hat{f}_{\text{FE}}$. Given observed transition samples $\{(s_{i},a_{i},s_{i+1})\}_{i=1}^{t-1}$ and the current state-action pair $(s_{t},a_{t})$, the function encoder predicts the next state $\hat{s}_{t+1}$ as: $\hat{s}_{t+1}=\hat{f}_{\text{FE}}(s_{t},a_{t})=\sum_{i=1}^{k}b_{i}\cdot g_{i}(s_{t},a_{t}),$ where $g_{i}(s_{t},a_{t})$ are pretrained basis functions, and $b_{i}$ are coefficients derived from a subset of transition samples. These coefficients $b_{i}$ additionally serve as a representation for $T_{\theta}$. Due to the properties of basis functions, these representations are fully informative and linear Ingebrand et al. (2024b). We concatenate them with the state to form an augmented input $\left(s_{t},b_{1},\ldots,b_{k}\right)$, which the policy uses as input. As the agent interacts with the environment, collecting new transitions $(s_{t},a_{t},s_{t+1})$, we refine the coefficients $b_{i}$ by solving Equation 5 with updated transition samples. Consequently, the agent receives an online representation of the dynamics. Note that with a fixed number of basis functions $k$, the computation, involving the inverse of a $k\times k$ matrix, remains efficient even for large samples. We denote the coefficients $(b_{1},\cdots,b_{k})$ as $b$.
@@ 4.3 Adaptive Shielding Mechanism
To ensure safety during policy execution, we propose an adaptive shielding mechanism that dynamically intervenes based on uncertainty in model predictions. This shield wraps any underlying policy $\pi$, adjusting actions to prevent unsafe outcomes. We illustrate the shielding process at timestep $t$. Given that the safety-regularized objective mitigates long-term cost violations, we adopt one-step prediction to minimize runtime overhead. For complex tasks, multi-step predictions can be employed.
We first introduce the necessary settings. The cost function is defined as $C(s_{t},a_{t},s_{t+1})=\mathbb{I}\left\{\nu(h(s_{t+1}),E_{t+1})\leq 0\right\}$, where $h:S\to\mathbb{R}^{n_{1}}$ extracts agent-centric safety features, $E_{t+1}\in\mathbb{R}^{n_{2}}$ captures environment features, and $\nu:\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}\to\mathbb{R}$ is Lipschitz continuous. By the Lipschitz property, the equation $\|h(s_{t+1})-h(s_{t})\|\leq\Delta_{\max}$ implies:
$\nu(h(s_{t+1}),E_{t+1})\geq\nu(h(s_{t}),E_{t})-L_{\nu}\Delta_{\max}.$ (8)
Thus, if $\nu(h(s_{t}),E_{t})>L_{\nu}\Delta_{\max}$, then $C(s_{t},a_{t},s_{t+1})=0$ for all $a_{t}\in A$. Since the value $\nu\left(h\left(s_{t}\right),E_{t}\right)$ can be computed at state $s_{t}$ before selecting action $a_{t}$ to assess its safety, we call it as the pre-safety indicator.
@@
    Action Generation: The policy $\pi$ generates $N$ candidate actions $\{a^{(i)}_{t}\}_{i=1}^{N}$ by sampling from its action distribution $\pi(\cdot\mid s_{t},b)$, where $b$ derived by a subset of transition samples up to time step $t$ explained in Section 4.2.
@@
    $\text{SafetyScore}(a^{(i)}_{t})=\nu\left(h_{t+1}(\hat{s}_{t+1}^{(i)}),\hat{E}_{t+1}\right)-2L_{\nu}\Gamma_{t},$ (10)
    where $\hat{E}_{t+1}$ represents predicted environment features and $\Gamma_{t}$ is the adaptive conformal prediction bound for $h_{t+1}(\hat{s}_{t+1}^{(i)})$ and $\hat{E}_{t+1}$ calibrated to maintain a $1-\delta$ safety probability. Actions are ranked by their safety scores, with positive scores indicating safety compliance.
@@
    Action Selection: Define the safe action set at state $s_{t}$ as $\hat{A}_{\text{safe}}(s_{t})=\{a^{(i)}_{t}:\text{SafetyScore}(a^{(i)}_{t})>0\}$ and sampled action set as $\hat{A}_{\text{sample}}=\{a^{(i)}_{t}\}_{i\in[N]}$ The shield executes the following selection rule:
    $a_{t}^{*}=\begin{cases}a\sim\mathcal{U}(\text{Top}_{k}(\hat{A}_{\text{safe}}(s_{t})),&\text{if }\hat{A}_{\text{safe}}\neq\emptyset,\\ \arg\max_{a\in\hat{A}_{\text{sample}}}\text{SafetyScore}(a),&\text{otherwise},\end{cases}$ (11)
@@
The following theorem demonstrates that an optimal policy, augmented with an adaptive shield, maximizes the expected cumulative discounted return while maintaining a tight bound on the average cost rate. See Appendix B for details.
@@
Given a Constrained Hidden Parameter MDP =$\mathcal{M}=\left\{S,A,\Theta,T,R,C,\gamma,P_{\Theta}\right\}$ with initial state $s_{0}\in S$ and failure probability $\delta\in(0,1)$, an optimal policy $\pi^{*}:S\times\Phi\rightarrow A$ augmented with an adaptive shield maximizes the expected cumulative discounted return $J_{R}(\pi^{*})$ while satisfying the average cost rate constraint: for $\theta\sim P_{\Theta}$ and some $0\leq\bar{\epsilon}\leq 1$,
$\phi^{\pi^{*}}\left(s,\theta\right)=\lim_{H\rightarrow\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\theta}}\left[\sum_{t=0}^{H-1}C(s_{t},a_{t},s_{t+1})\mid s_{0}=s,\theta\right]\leq\delta+\bar{\epsilon}(1-\delta).$ (12)
Given this bound, if safe actions exist at each step, this theorem proves that our algorithm achieves a low average cost rate constraint, governed by the ACP failure probability, i.e., $\phi^{\pi^{*}}\left(s,\theta\right)\leq\delta$.
@@
The ACP provides a probabilistic guarantee on the deviation between the predicted state $\hat{s}_{t+1}=$ $\hat{f}\left(s_{t},a_{t}\right)$ and the true state $s_{t+1}$: $\left(\left\|s_{t+1}-\hat{s}_{t+1}\right\|\leq\Gamma_{t}\right)\geq 1-\delta,$ where $\Gamma_{t}$ is the confidence region at time $t+1$. Since $\nu$ is Lipschitz continuous with constant $L_{\nu}$, we bound the difference in the safety margin between the true and predicted states: $\nu\left(h\left(s_{t+1}\right),E_{t+1}\right)\geq\nu(h\left(\hat{s}_{t+1}\right),\hat{E}_{t+1})-L_{\nu}\left\|h\left(s_{t+1}\right)-h\left(\hat{s}_{t+1}\right)\right\|-L_{\nu}\left\|E_{t+1}-\hat{E}_{t+1}\right\|.$ We assume that prediction errors are bounded by accurate next-state predictions: $\|h\left(s_{t+1}\right)-h\left(\hat{s}_{t+1}\right)\|\leq\Gamma_{t}$ and $\left\|E_{t+1}-\hat{E}_{t+1}\right\|\leq\Gamma_{t}$. The set of safe actions is defined based on the predicted margin: $\hat{A}_{\mathrm{safe}}\left(s_{t}\right)=\{a\in A\mid\nu(h(\hat{f}\left(s_{t},a\right)),\hat{E}_{t+1})>2L_{\nu}\Gamma_{t}\}.$ If an action is selected from $\hat{A}_{\text{safe }}$, we guarantee $\nu\left(h\left(s_{t+1}\right),E_{t+1}\right)>0$, ensuring a safe state at $t+1$ The final bound depends on the failure probability of state prediction and the probability of selecting safe actions. ∎

### v2

@@ 4 Approach
Our approach has three main components. First, we introduce a safety-regularized optimization (SRO). Through this optimization, the policy is trained to converge toward a zero-violation policy. Second, we use a function encoder to represent underlying dynamics $T_{\phi}$, enabling online adaptation. Finally, we use this dynamics representation to construct an adaptive shield. The shield adjusts safe regions by conformal prediction and blocks unsafe actions online.
4.1 Safety-Regularized Optimization
@@
To promote safe policy learning, we introduce a safety measure, $Q^{\pi}_{\text{safe}}(s,a,b_{\phi})$, which defines a local cost-sensitivity score for an action $a\sim\pi(\cdot|s,b_{\phi})$, given the learned representation $b_{\phi}\in\mathbb{R}^{k}$. The score measures how strongly the expected cumulative cost varies in a neighborhood of $a$, weighted by the policy’s local action density. Higher values of $Q^{\pi}_{\text{safe}}$ indicate actions with lower long-term costs under policy $\pi$. Since we aim to minimize the cost action-value function $Q^{\pi}_{C}$, higher $Q^{\pi}_{\text{safe}}$ values correspond to lower $Q^{\pi}_{C}$ values. Based on this intuition, we define $Q^{\pi}_{\text{safe}}(s,a,b_{\phi})$ for an action $a\sim\pi(\cdot|s,b_{\phi})$ as:
$Q^{\pi}_{\text{safe}}(s,a,b_{\phi})=-\frac{\int_{B(a,\epsilon)\cap A}\pi(x\mid s,b_{\phi})Q^{\pi}_{C}(s,x,b_{\phi})dx}{V^{\pi}_{C}(s,b_{\phi})+\epsilon}$ (6)
@@
where $\epsilon>0$ is a small constant ensuring numerical stability, and $B(a,\epsilon)$ denotes a small ball of radius $\epsilon$ centered at $a$.
This formulation bounds the value in (1,0](-1,0] by its design. For continuous action spaces, the probability of taking a specific action is always 0, so we integrate the value over a small interval including that action. For practical implementation, we use a Monte Carlo approximation to the integral by sampling several values around an action $a$ and then aggregating them. See Appendix C for details.
We interpret $Q^{\pi}_{\text{safe}}(s,a,b_{\phi})$ as a scalar value during optimization (i.e., we apply a stop-gradient to this term). This ensures that the policy updates focus on shifting probability mass toward safer regions. The resulting gradient for the safety objective $J_{\text{safe}}(\pi)={\mathbb{E}}[Q^{\pi}_{\text{safe}}(s,a,b_{\phi})]$ is approximated as:
@@
$\nabla_{\theta}J_{\text{safe}}(\pi)\approx\mathbb{E}\left[\nabla_{\theta}\log\pi_{\theta}\left(a\mid s,b_{\phi}\right)Q_{\text{safe}}^{\pi}\left(s,a,b_{\phi}\right)\right].$ (7)
A value of $Q^{\pi}_{\text{safe}}(s,a,b_{\phi})$ near 0 implies that $\|\nabla_{\theta}J_{\text{safe}}(\pi)\|$ is small, so the regularizer contributes little to the policy update. This occurs when either (i) the predicted cumulative cost in a local neighborhood of $a$ is small, or (ii) the policy assigns negligible probability mass to that neighborhood, leading to a small numerator in (6). In contrast, values near 1-1 indicate that regions of high policy probability are associated with large long-term cost, producing a large gradient magnitude that discourages the policy from concentrating on locally high-risk action regions. See Appendix J for details on the design choice.
To integrate this safety measure into policy optimization, we define an augmented action-value function, $Q^{\pi}_{\text{aug}}(s,a,b_{\phi})=Q_{R}^{\pi}(s,a,b_{\phi})+\alpha Q_{\text{safe}}^{\pi}(s,a,b_{\phi})$, where $Q_{R}^{\pi}(s,a,b_{\phi})$ is the reward action-value function, and $\alpha\geq 0$ is a hyperparameter balancing safety and reward. Our safety-regularized optimization (SRO) updates the policy by maximizing the corresponding augmented objective:
$J_{\text{aug}}(\pi)=\mathbb{E}_{\phi\sim P_{\phi},\,s_{0}\sim\mu_{0}(\cdot\mid\phi),\,a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q^{\pi}_{\text{aug}}(s_{0},a_{0},b_{\phi})\right].$ (8)
@@
A larger $\alpha$ encourages the policy to prioritize safe actions that result in low long-term costs or to select under-explored actions with lower assigned probabilities. Next, we introduce a proposition which justifies this choice of objective.
Proposition 1 (Reward Consistency within Zero-Violation Policies).
@@
Let $\Pi_{\text{zero-violation}}$ denote the set of zero-violation policies, defined as $\{\pi\mid J_{C}(\pi)=0\}$. Then, for any $\alpha\geq 0$, the optimal policy obtained by maximizing the augmented objective function $J_{\text{aug}}(\pi)$ within $\Pi_{\text{zero-violation}}$ is equivalent to the optimal policy obtained by maximizing the standard reward objective $J_{R}(\pi)$ within the same set of policies.
@@
For any policy within the zero-violation set, all actions sampled from the policy lead to $Q_{C}^{\pi}(s,a,b_{\phi})=0$. By our design of the safety term, this condition implies $Q_{\text{safe}}^{\pi}(s,a,b_{\phi})=0$. Substituting this into our augmented objective, $J_{\text{aug}}(\pi)$ simplifies to $J_{R}(\pi)$. ∎
Proposition 1 demonstrates that SRO does not degrade performance unnecessarily when an agent already behaves safely. Specifically, it guarantees that if we focus only on the set of policies that satisfy all safety constraints, maximizing the safety-regularized objective is equivalent to maximizing the standard reward objective. SRO preserves the standard monotonic improvement and worst-case constraint bounds of TRPO and CPO, see Appendix B for details.
@@ 4.2 Inferring Hidden Parameters Online
To infer the underlying dynamics $T_{\phi}$ and predict the next state $s_{t+1}$ based on transition samples and $(s_{t},a_{t})$, we use a function encoder, denoted by $\hat{f}_{\text{FE}}$. Given observed transition samples $\{(s_{i},a_{i},s_{i+1})\}_{i=1}^{t-1}$ and the current state-action pair $(s_{t},a_{t})$, the function encoder predicts the next state $\hat{s}_{t+1}$ as: $\hat{s}_{t+1}=\hat{f}_{\text{FE}}(s_{t},a_{t})=\sum_{i=1}^{k}b_{i}\cdot g_{i}(s_{t},a_{t}),$
where $g_{i}(s_{t},a_{t})$ are pretrained basis functions, and $b_{i}$ are coefficients derived from a subset of transition samples. These coefficients $b_{i}$ additionally serve as a representation for $T_{\phi}$. Due to the properties of basis functions, these coefficients provide a linear, low-dimensional representation of the underlying dynamics (tyler2024zero_shot). We concatenate them with the state to form an augmented input $\left(s_{t},b_{1},\ldots,b_{k}\right)$, which the policy uses as input. As the agent interacts with the environment, collecting new transitions $(s_{t},a_{t},s_{t+1})$, we refine the coefficients $b_{i}$ by solving Equation 5 with updated transition samples. Consequently, the agent receives an online representation of the dynamics. Note that with a fixed number of basis functions $k$, the computation, involving the inverse of a $k\times k$ matrix, remains efficient even for large samples. We denote the coefficients $(b_{1},\cdots,b_{k})$ as $b_{\phi}$ for the dynamics $T_{\phi}$. We interpret this procedure as a lightweight online system identification, where short transition histories are mapped to a low-dimensional representation of the environment dynamics without retraining the policy.
@@ 4.3 Adaptive Shielding Mechanism
To ensure safety during policy execution, we propose an adaptive shielding mechanism that dynamically intervenes based on uncertainty in model predictions. This shield wraps any underlying policy $\pi$, adjusting actions to prevent unsafe outcomes. We illustrate the shielding process at timestep $t$.
We first introduce the necessary settings. The cost function is defined using an indicator function $\mathbb{I}$ as $C(s_{t},a_{t},s_{t+1})=\mathbb{I}\left\{\nu(e(s_{t+1}),E_{t+1})\leq 0\right\}$, where $e:S\to\mathbb{R}^{n_{1}}$ extracts agent-centric safety features, $E_{t+1}\in\mathbb{R}^{n_{2}}$ captures environment features, and $\nu:\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}\to\mathbb{R}$ is Lipschitz continuous with Lipschitz constant $L_{\nu}$. We assume the agent-centric safety features change smoothly, i.e., $\left\|e\left(s_{t+1}\right)-e\left(s_{t}\right)\right\|\leq\Delta_{\max},$ where $\Delta_{\max}$ is a bound on the per-step feature change. By the Lipschitz property, the equation $\|e(s_{t+1})-e(s_{t})\|\leq\Delta_{\max}$ implies:
$\nu(e(s_{t+1}),E_{t+1})\geq\nu(e(s_{t}),E_{t})-L_{\nu}\Delta_{\max}.$ (9)
Thus, if $\nu(e(s_{t}),E_{t})>L_{\nu}\Delta_{\max}$, then $C(s_{t},a_{t},s_{t+1})=0$ for all $a_{t}\in A$. Since the value $\nu\left(e\left(s_{t}\right),E_{t}\right)$ can be computed at state $s_{t}$ before selecting action $a_{t}$ to assess its safety, we call it as the pre-safety indicator.
@@
    Action Generation: The policy $\pi$ generates $N$ candidate actions $\{a^{(i)}_{t}\}_{i=1}^{N}$ by sampling from its action distribution $\pi(\cdot\mid s_{t},b_{\phi})$, where $b_{\phi}$ derived by a subset of transition samples up to time step $t$ explained in Section 4.2.
@@
    $\text{SafetyScore}(a^{(i)}_{t})=\nu\left(e(\hat{s}_{t+1}^{(i)}),\hat{E}_{t+1}\right)-2L_{\nu}\Gamma_{t},$ (11)
    where $\hat{E}_{t+1}$ represents predicted environment features and $\Gamma_{t}$ is the adaptive conformal prediction bound for $\hat{s}_{t+1}^{(i)}$ and $\hat{E}_{t+1}$ calibrated to target an empirical miscoverage level $\delta$ for the next-state prediction error. Actions are ranked by their safety scores, with positive scores indicating safety compliance. We treat $\Gamma_{t}$ as an adaptive uncertainty estimate to recent observations, showing a time-varying uncertainty estimate.
@@
    Action Selection: Define the safe action set at state $s_{t}$ as $\hat{A}_{\text{safe}}(s_{t})=\{a^{(i)}_{t}:\text{SafetyScore}(a^{(i)}_{t})>0\}$ and sampled action set as $\hat{A}_{\text{sample}}=\{a^{(i)}_{t}\}_{i\in[N]}$ . The shield executes the following selection rule:
    $a_{t}^{*}=\begin{cases}a\sim\mathcal{U}(\text{Top}_{k}(\hat{A}_{\text{safe}}(s_{t}))),&\text{if }\hat{A}_{\text{safe}}\neq\emptyset,\\ \arg\max_{a\in\hat{A}_{\text{sample}}}\text{SafetyScore}(a),&\text{otherwise},\end{cases}$ (12)
@@
When the shield predicts multiple steps $h$ ahead, we repeat the procedure for steps 2, 3, 4, aggregating the safety score over future steps. However, long-term predictions often increase compounding errors and runtime. Thus, we typically use a shorter prediction horizon such as $h=1$ or $h=2$.
The following theorem characterizes how prediction errors in the learned dynamics model connects with a bound on the average cost rate of a policy augmented with the adaptive shield. See Appendix A for details.
@@
Given a Constrained Hidden Parameter MDP $\mathcal{M}=(S,A,\Phi,T,R,C,\gamma,P_{\Phi})$ with initial state $s_{0}\in S$ and failure probability $\delta\in(0,1)$, assume the prediction error of the learned dynamics model satisfies $\mathbb{P}(\|s_{t+1}-\hat{s}_{t+1}\|\leq\Gamma_{t})\geq 1-\delta$. Then, for any policy $\pi$ augmented with the adaptive shield and $\phi\sim P_{\Phi}$, there exists $0\leq\bar{\epsilon}\leq 1$ such that the average cost rate constraint: for $\phi\sim P_{\Phi}$ and some $0\leq\bar{\epsilon}\leq 1$,
$\xi^{\pi}\left(s,\phi\right)$ $=\lim_{H\rightarrow\infty}\frac{1}{H}\mathbb{E}_{\pi,T_{\phi}}\left[\sum_{t=0}^{H-1}C(s_{t},a_{t},s_{t+1})\mid s_{0}=s,\phi\right]$
$\leq\delta+\bar{\epsilon}(1-\delta).$
Given this bound, if safe actions exist at each step, this theorem proves that our algorithm achieves a low average cost rate constraint, governed by the ACP failure probability, i.e., $\xi^{\pi^{*}}\left(s,\phi\right)\leq\delta$.
@@
[…]
Guaranteeing $\bar{\epsilon}=0$ requires complete knowledge of dynamics, which is impossible in many settings. Instead, the theorem distinguishes the uncertainty we can reduce through learning from the uncontrollable risks of environments.